Find the wavelength of the spectral line that corresponds to a transition in hydrogen from the n = 10 state to the ground state. In what part of the spectrum is this?


1
Expert's answer
2021-11-25T16:23:58-0500

Given:

"n=10,\quad m=1"


The Rydberg formula says

"\frac{1}{\lambda}=R\left(\frac{1}{m^2}-\frac{1}{n^2}\right)"

"\frac{1}{\lambda}=1.09677×10^7\:\rm m^{−1}\left(\frac{1}{1^2}-\frac{1}{10^2}\right)"

"\frac{1}{\lambda}=1.086\times 10^7\:\rm m^{-1}"

"\lambda=92\times 10^{-9}\;\rm m=92\: nm"

This is the Lyman series.